# Exploring Hubble Tension Alleviation through Neutrino-Coupled Perturbed $f(R)$ Gravity

### v1

@@ Abstract
I Introduction
One of the most pressing challenges in modern cosmology is the persistent discrepancy in determining the Hubble constant, $H_{0}$. This so-called ”Hubble tension” has become a major enigma, drawing extensive attention from the scientific community. The disagreement primarily arises from two independent methodologies, which consistently yield conflicting estimates of $H_{0}$.
The first approach is based on observations of the cosmic microwave background (CMB), as analyzed by the Planck collaboration over a seven-year period Planck13 ; Planck15 ; Planck18 . The most recent estimate from Planck 2018 places the Hubble constant at $H_{0}=67.4\pm 0.5,\text{km s}^{-1}\text{Mpc}^{-1}$ Planck18 . In contrast, the second method, employed by the SH0ES collaboration using data from the Hubble Space Telescope (HST), provides a significantly higher estimate, with the latest value reported as $H_{0}=74.03\pm 1.42,\text{km s}^{-1}\text{Mpc}^{-1}$ HST18 ; HST19 . The current discrepancy between these measurements exceeds $4.4\sigma$, and the tension has only intensified in recent years.
Several independent measurements of the Hubble constant ($H_{0}$) have been reported in the literature. These values are obtained using different observational techniques, such as the cosmic microwave background (CMB), Type Ia supernovae (SNe Ia), gravitational lensing, and standard candles like the Tip of the Red Giant Branch (TRGB) and Miras. The following table summarizes the reported values:
Study $H_{0}$ (km s-1 Mpc-1) Method
Planck 2018 Planck18 $67.4\pm 0.5$ CMB (Planck)
SH0ES (HST) HST19 $74.03\pm 1.42$ SNe Ia (Cepheid Calibration)
CCHP (Carnegie-Chicago Hubble Program) FreedmanTRGB20 $69.6\pm 0.8\pm 1.7$ TRGB (Tip of the Red Giant Branch)
H0LiCOW Wonglens19 $73.3^{+1.7}_{-1.8}$ Gravitational Lensing (Time Delay)
H0LiCOW Weilens20 $75.3^{+3.0}_{-2.9}$ Gravitational Lensing (Updated Analysis)
Baxter et al. (2020) BaxterCMBlens20 $73.5\pm 5.3$ CMB Lensing
Huang et al. (2019) HuangMiras19 $72.7\pm 4.6$ Miras (Variable Stars)
Table 1: Summary of recent $H_{0}$ measurements from different observational methods.
The tension between the early universe (Planck CMB) and late universe (SNe Ia, gravitational lensing, etc.) measurements remains a significant challenge in modern cosmology.
This discrepancy can be interpreted as a mismatch between early and late Universe observations VerdeTR19 . The SH0ES collaboration utilizes data from local distance indicators, while the Planck collaboration derives its estimate from observations spanning a wide redshift range ($11000<z<$ 1100) within the $\Lambda\text{CDM}$ model. Resolving this issue may require more than reconciling observational discrepancies—it could necessitate revisiting the theoretical framework of cosmology and considering modifications to the standard model.
The standard Cold Dark Matter (CDM) model, while providing a compelling explanation for many cosmological phenomena, faces several challenges, including the cosmic coincidence and fine-tuning problems. Alternative approaches to addressing these issues generally fall into two broad categories. The first class of models focuses on modifying the properties of the matter field, leading to dynamic dark energy models such as quintessence agr-cosmo-de-quintessence-01 , tachyon agr-cosmo-de-tachyon-01 , k-essence agr-cosmo-de-k-essence-01 , phantom agr-cosmo-de-phantom-01 , Chaplygin gas agr-cosmo-de-chaplygin-01 , holographic dark energy agr-cosmo-de-holo-01 ; agr-cosmo-de-holo-02 ; agr-cosmo-de-holo-04 , and agegraphic dark energy agr-cosmo-de-agegraphic-01 ; agr-cosmo-de-agegraphic-02 , among others.
@@ V Numerical Analysis
To assess the viability and effectiveness of the model under investigation, we conduct a series of Markov Chain Monte Carlo (MCMC) simulations using the publicly available cosmological parameter estimation code, MontePython-v3.222https://github.com/brinckmann/montepython_public 54 ; 55 . This software is interfaced with our modified version of the Boltzmann solver CLASS 56 ; 57 , which incorporates the necessary theoretical modifications relevant to our study.
To ensure the robustness of the MCMC sampling and verify that the generated Markov chains have reached a statistically reliable equilibrium, we impose the Gelman-Rubin convergence criterion 58 , requiring that the deviation from unity satisfies $|R-1|\lesssim 0.01$. This criterion ensures that different chains, initialized with independent random starting points, converge to the same posterior distribution, indicating a sufficient exploration of the parameter space.
After completing the MCMC runs, we employ the GetDist analysis package 59 to process the sampled chains. This involves computing marginalized posterior distributions, estimating best-fit values and confidence intervals for the cosmological parameters, and generating visualizations such as one- and two-dimensional contour plots. These post-processing steps allow us to extract meaningful physical constraints and compare our model’s predictions against observational data. Here’s an improved version of your text with more clarity, completeness, and precision:
@@
We utilize the Pantheon+ compilation, which consists of 896 SNe Ia spanning the redshift range $2.30.12<z<$ 2.3 42 . This dataset provides precise luminosity distance measurements and serves as a crucial probe for constraining the late-time expansion history of the universe salehi et al. (2020).
@@
We include 32 cosmic chronometer (CC) measurements of the Hubble parameter $H(z)$ spanning the redshift range $1.9650.07<z<$ 1.965, as listed in Table 3. The CC technique provides model-independent constraints on $H(z)$ by using the differential age method to determine the expansion rate of the universe at different epochs.
@@ VI Data Integration and Likelihood Functions in MontePython
Cosmic Microwave Background (CMB)
MontePython constrains cosmological models using CMB data by comparing theoretical and observed power spectra. The CMB likelihood function is:
$\mathcal{L}_{\text{CMB}}\propto\exp\left(-\frac{1}{2}\left[\textbf{d}_{\text{obs}}-\textbf{d}_{\text{theory}}\right]^{T}\textbf{C}^{-1}\left[\textbf{d}_{\text{obs}}-\textbf{d}_{\text{theory}}\right]\right),$ (45)
where $\textbf{d}_{\text{obs}}$ and $\textbf{d}_{\text{theory}}$ are the observed and theoretical power spectra, and C is the covariance matrix.
@@
CC data provides expansion rate measurements, incorporated using the likelihood function:
$\mathcal{L}_{\text{CC}}\propto\exp\left(-\frac{1}{2}\sum_{i}\left[\frac{H(z_{i})_{\text{model}}-H(z_{i})_{\text{data}}}{\sigma_{H}(z_{i})}\right]^{2}\right).$ (46)
Here, $H(z_{i})_{\text{model}}$ and $H(z_{i})_{\text{data}}$ are the theoretical and observed expansion rates, respectively, with $\sigma_{H}(z_{i})$ as uncertainty.
@@
BAO data constrains cosmology through the BAO scale, with likelihood:
$\mathcal{L}_{\text{BAO}}\propto\exp\left(-\frac{1}{2}\frac{\left(D_{V}^{\text{model}}-D_{V}^{\text{data}}\right)^{2}}{\sigma_{D_{V}}^{2}}\right),$ (47)
where $D_{V}^{\text{model}}$ and $D_{V}^{\text{data}}$ are theoretical and observed BAO scales, and $\sigma_{D_{V}}$ is the uncertainty.
Pantheon Supernovae (SN)
Pantheon data compares theoretical and observed distance moduli using:
$\mathcal{L}_{\text{SN}}\propto\exp\left(-\frac{1}{2}\sum_{j}\left[\frac{\mu_{j}^{\text{model}}-\mu_{j}^{\text{data}}}{\sigma_{\mu}(j)}\right]^{2}\right),$ (48)
where $\mu_{j}^{\text{model}}$ and $\mu_{j}^{\text{data}}$ are the predicted and observed distance moduli, with $\sigma_{\mu}(j)$ as the uncertainty.
Lensing Data
MontePython evaluates lensing power spectra through the likelihood function:
$\mathcal{L}_{\text{lensing}}$ $\propto\exp\left(-\frac{1}{2}\left[\textbf{d}_{\text{lensing,obs}}-\textbf{d}_{\text{lensing,theory}}\right]^{T}\right.$
×$\quad\times\left.\textbf{C}_{\text{lensing}}^{-1}\left[\textbf{d}_{\text{lensing,obs}}-\textbf{d}_{\text{lensing,theory}}\right]\right),$ (49)
where $\textbf{d}_{\text{lensing,obs}}$ and $\textbf{d}_{\text{lensing,theory}}$ are observed and theoretical lensing power spectra, and $\textbf{C}_{\text{lensing}}$ is the covariance matrix.
MCMC Sampling with MontePython
MontePython combines all likelihoods:
where $\theta$ represents cosmological parameters. The posterior probability is:
$\mathcal{P}(\theta|\text{data})\propto\mathcal{L}(\text{data}|\theta)\cdot\text{Prior}(\theta).$ (51)
MontePython employs Markov Chain Monte Carlo (MCMC) sampling to explore parameter space and derive posterior distributions. Table IV shows the flat priors for the cosmological parameters based on (Ratra1 ; Aghanim et al. (2018); Aghanim et al. (2018); Di Valentino et al. (2021); Cardona & Sabogal (2023))
@@
$\Omega_{b}h^{2}$ $[0.00,0.10]$
$\Omega_{c}h^{2}$ $[0.00,0.15]$
$\tau$ $[0.00,0.80]$
$n_{s}$ $[0.80,1.20]$
$\log[10^{10}A_{s}]$ $[1.60,3.90]$
$100\theta_{MC}$ $[0.50,10.50]$
${\Omega_{\nu}}$ $[0.00,0.009]$
@@ VIII Hubble Tension analysis
${\Omega_{b}h^{2}}$ $0.02226\pm 0.00022$ $0.02228\pm 0.00023$ $0.02234\pm 0.00020$ $0.02236\pm 0.00021$ $0.02219\pm 0.00018$
${\Omega_{c}h^{2}}$ $0.1196\pm 0.0037$ $0.1194\pm 0.0041$ $0.1191\pm 0.0035$ $0.1194\pm 0.0035$ $0.1187\pm 0.0025$
${100\theta_{MC}}$ $1.0410\pm 0.0006$ $1.0408\pm 0.0007$ $1.0411\pm 0.0006$ $1.04115\pm 0.0005$ $1.0408\pm 0.0005$
${\tau}$ $0.059\pm 0.008$ $0.061\pm 0.009$ $0.0551\pm 0.008$ $0.055\pm 0.008$ $0.054^{+0.006}_{-0.007}$
${\mathrm{ln}(10^{10}A_{s})}$ $3.047\pm 0.02$ $3.045\pm 0.017$ $3.043\pm 0.019$ $3.046\pm 0.018$ $3.042^{+0.016}_{-0.017}$
${n_{s}}$ $0.971\pm 0.009$ $0.970\pm 0.009$ $0.974\pm 0.008$ $0.964\pm 0.008$ $0.969\pm 0.007$
@@
CMB + Lensing $69.6\pm 2.8$ $0.61\sigma$ $1.26\sigma$
CMB + BAO + Lensing $70.15\pm 1.6$ $1.23\sigma$ $1.18\sigma$
CMB + Lensing + CC $70.52\pm 1.92$ $1.15\sigma$ $1.22\sigma$
CMB + Lensing + Pantheon $70.88\pm 1.94$ $1.22\sigma$ $1.10\sigma$
CMB + All Data (BAO, CC, Pantheon, Lensing) $70.18\pm 1.79$ $1.51\sigma$ $1.44\sigma$
@@
${\Omega_{b}h^{2}}$ $0.02230\pm 0.00022$ $0.02229\pm 0.00024$ $0.02232\pm 0.00021$ $0.02233\pm 0.00023$ $0.02210\pm 0.00016$
${\Omega_{c}h^{2}}$ $0.1198\pm 0.0036$ $0.1194\pm 0.0039$ $0.1195\pm 0.0033$ $0.1197\pm 0.0034$ $0.1186\pm 0.0022$
${100\theta_{MC}}$ $1.0408\pm 0.0006$ $1.0409\pm 0.0007$ $1.0410\pm 0.0005$ $1.04111\pm 0.0006$ $1.0406\pm 0.0004$
${\tau}$ $0.058\pm 0.008$ $0.059\pm 0.009$ $0.0550\pm 0.008$ $0.055\pm 0.009$ $0.053^{+0.005}_{-0.007}$
${\mathrm{ln}(10^{10}A_{s})}$ $3.048\pm 0.019$ $3.044\pm 0.018$ $3.045\pm 0.017$ $3.045\pm 0.017$ $3.041^{+0.015}_{-0.016}$
${n_{s}}$ $0.973\pm 0.009$ $0.971\pm 0.010$ $0.972\pm 0.008$ $0.966\pm 0.009$ $0.968\pm 0.006$
@@
CMB + Lensing $69.82\pm 3.12$ $0.74\sigma$ $0.95\sigma$
CMB + BAO + Lensing $70.37\pm 2.32$ $1.14\sigma$ $1.03\sigma$
CMB + Lensing + CC $70.26\pm 2.38$ $1.13\sigma$ $0.98\sigma$
CMB + Lensing + Pantheon $70.57\pm 2.54$ $1.17\sigma$ $0.88\sigma$
CMB + All Data (BAO, CC, Pantheon, Lensing) $70.46\pm 2.01$ $1.42\sigma$ $1.12\sigma$

### v2

@@ Abstract
We investigate the Hubble constant ($H_{0}$) and $S_{8}$ tensions within the framework of perturbed $f(R)$ gravity models, both with and without neutrino coupling. By analyzing a comprehensive set of cosmological data, including cosmic microwave background (CMB), baryon acoustic oscillations (BAO), cosmic chronometers (CC), Pantheon supernovae, and lensing, we assess the capability of these models to mitigate discrepancies between early- and late-time measurements. The perturbed $f(R)$ model without neutrino coupling demonstrates moderate improvements in inferred $H_{0}$ values, with tensions ranging from $0.61\sigma$ to $1.51\sigma$ relative to Planck 2018 and from $1.10\sigma$ to $1.44\sigma$ relative to R22. However, the coupled model incorporating neutrino interactions exhibits greater potential to alleviate the Hubble tension, reducing the discrepancy with Planck 2018 to $0.74\sigma$ and with R22 to $0.95\sigma$. Similarly, the coupled model improves consistency in the $S_{8}$ parameter, reducing the tension with Planck to $1.13\sigma$ and lowering tensions with KiDS-1000 and DES-Y3 to $0.70\sigma$ and $0.23\sigma$, respectively. These results highlight the enhanced performance of the neutrino-coupled perturbed $f(R)$ model in addressing cosmological tensions. While not resolving all discrepancies, the improved agreement with late-time measurements underscores the potential of modified gravity and neutrino interactions to account for deviations in cosmological observations.
I Introduction
The Hubble constant ($H_{0}$) represents the current expansion rate of the Universe and remains one of the most contentious parameters in modern cosmology due to the persistent discrepancy between measurements from early- and late-Universe probes. The two most widely cited methods for determining $H_{0}$ are based on observations of the cosmic microwave background (CMB) and Type Ia supernovae (SNe Ia).
The first approach relies on CMB observations, which provide a precise estimate of the Hubble constant by analyzing the angular size of the acoustic peaks in the CMB power spectrum. The Planck collaboration, using seven years of data, reported the most recent estimate in 2018 as $H_{0}=67.4\pm 0.5\,\text{km s}^{-1}\text{Mpc}^{-1}$ Planck18 . This value is consistent with predictions from the standard $\Lambda$CDM model but is in significant tension with measurements from late-Universe probes.
In contrast, the second approach, employed by the SH0ES (Supernovae H0 for the Equation of State) collaboration, is based on the Cepheid-calibrated distance ladder using Type Ia supernovae data from the Hubble Space Telescope (HST). The most recent SH0ES result reports $H_{0}=74.03\pm 1.42\,\text{km s}^{-1}\text{Mpc}^{-1}$ HST18 ; HST19 , which exceeds the Planck value by more than $4.4\sigma$. This significant tension between early- and late-Universe measurements has only intensified in recent years.
Independent measurements using alternative methods have yielded intermediate or conflicting results. For instance, the Carnegie-Chicago Hubble Program (CCHP) reported a value of $H_{0}=69.6\pm 0.8\pm 1.7\,\text{km s}^{-1}\text{Mpc}^{-1}$ based on the Tip of the Red Giant Branch (TRGB) method FreedmanTRGB20 , which provides a distance calibration independent of Cepheids. Similarly, the H0LiCOW collaboration, using gravitational lensing time delays, reported two separate measurements: $H_{0}=73.3^{+1.7}_{-1.8}\,\text{km s}^{-1}\text{Mpc}^{-1}$ Wonglens19 and an updated value of $H_{0}=75.3^{+3.0}_{-2.9}\,\text{km s}^{-1}\text{Mpc}^{-1}$ Weilens20 .
Additional measurements from CMB lensing, variable stars, and other astrophysical probes further complicate the picture. Baxter et al. (2020) obtained a value of $H_{0}=73.5\pm 5.3\,\text{km s}^{-1}\text{Mpc}^{-1}$ from CMB lensing BaxterCMBlens20 , while Huang et al. (2019) reported $H_{0}=72.7\pm 4.6\,\text{km s}^{-1}\text{Mpc}^{-1}$ using Miras (variable stars) HuangMiras19 . These results underscore the ongoing difficulty in reconciling early- and late-Universe estimates of the Hubble constant, which remains one of the most significant challenges in modern cosmology. Addressing this discrepancy requires exploring new physics beyond the standard cosmological model, including modifications to gravity, dark energy models, and possible systematic errors in observational data. This discrepancy can be interpreted as a mismatch between early and late Universe observations VerdeTR19 . Another significant tension arises from measurements of the $\sigma_{8}$ parameter, which reflects the amplitude of matter density fluctuations on scales of $8h^{-1}$ Mpc, where $h$ is the reduced Hubble constant. A related parameter, $S_{8}$, is commonly used to quantify the strength of matter density fluctuations and is defined as: $S_{8}=\sigma_{8}\sqrt{\frac{\Omega_{m}}{0.3}}$, where $\Omega_{m}$ is the matter density parameter. The parameter $S_{8}$ provides a useful measure of the amplitude of density fluctuations in the universe and plays a crucial role in understanding cosmological structure formation. Matter density fluctuations seed the formation of cosmic structures, such as galaxies and galaxy clusters. By comparing measurements of large-scale structures from galaxy surveys and cosmic microwave background (CMB) data with theoretical predictions, $S_{8}$ serves as a key parameter for constraining cosmological models.
@@ I Introduction
The standard $\Lambda$ Cold Dark Matter ($\Lambda$CDM) model, while providing a compelling explanation for many cosmological phenomena, faces several challenges, including the cosmic coincidence and fine-tuning problems. Alternative approaches to addressing these issues generally fall into two broad categories. The first class of models focuses on modifying the properties of the matter field, leading to dynamic dark energy models such as quintessence agr-cosmo-de-quintessence-01 , tachyon agr-cosmo-de-tachyon-01 , k-essence agr-cosmo-de-k-essence-01 , phantom agr-cosmo-de-phantom-01 , Chaplygin gas agr-cosmo-de-chaplygin-01 , holographic dark energy agr-cosmo-de-holo-01 ; agr-cosmo-de-holo-02 ; agr-cosmo-de-holo-04 , and agegraphic dark energy agr-cosmo-de-agegraphic-01 ; agr-cosmo-de-agegraphic-02 , among others.
@@ V Numerical Analysis
We utilize the Pantheon+ compilation, which consists of 896 SNe Ia spanning the redshift range $0.12<z<2.3$ 42 . This dataset provides precise luminosity distance measurements and serves as a crucial probe for constraining the late-time expansion history of the universe salehi et al. (2020).
@@
[…]
@@ VI Data Integration and Likelihood Functions in MontePython
MontePython (Audren2013MontePython, ; Brinckmann2019MontePython3, ) is a flexible cosmological inference framework that interfaces with the Boltzmann code CLASS (Blas2011CLASS, ) and supports a wide range of cosmological and astrophysical likelihoods. It enables robust parameter estimation through Bayesian inference using Markov Chain Monte Carlo (MCMC) or nested sampling techniques.
The total likelihood for a given parameter set $\theta$ is constructed as the product of individual likelihoods:
with each term representing the likelihood contribution from an independent cosmological probe.
CMB Likelihoods
MontePython incorporates official likelihoods from Planck 2018 (Planck2018, ). The likelihood is constructed from differences between observed and theoretical CMB temperature and polarization power spectra:
$\mathcal{L}_{\text{CMB}}\propto\exp\left(-\frac{1}{2}\Delta\mathbf{C}_{\ell}^{\mathrm{T}}\mathbf{C}_{\text{CMB}}^{-1}\Delta\mathbf{C}_{\ell}\right),$ (46)
where $\Delta\mathbf{C}_{\ell}=\mathbf{C}_{\ell}^{\text{obs}}-\mathbf{C}_{\ell}^{\text{theory}}$, and $\mathbf{C}_{\text{CMB}}$ is the covariance matrix for the angular power spectra.
@@
The CC likelihood utilizes Hubble parameter measurements $H(z)$ derived from differential aging of passively evolving galaxies. The likelihood function takes the Gaussian form:
$\mathcal{L}_{\text{CC}}\propto\exp\left(-\frac{1}{2}\Delta\mathbf{H}^{\mathrm{T}}\mathbf{C}_{\text{CC}}^{-1}\Delta\mathbf{H}\right),$ (47)
where $\Delta\mathbf{H}=\mathbf{H}_{\text{obs}}-\mathbf{H}_{\text{model}}$, and $\mathbf{C}_{\text{CC}}$ is the diagonal or full covariance matrix depending on the dataset used. Integration of CC data into MontePython is straightforward via the data/ likelihood module.
@@
BAO data constrain cosmological distances via features imprinted in the galaxy correlation function. MontePython includes likelihoods from 6dFGS, SDSS MGS, BOSS DR12, and eBOSS. The likelihood compares observed and theoretical distance combinations:
$\mathcal{L}_{\text{BAO}}\propto\exp\left(-\frac{1}{2}\Delta\mathbf{D}^{\mathrm{T}}\mathbf{C}_{\text{BAO}}^{-1}\Delta\mathbf{D}\right),$ (48)
where $\Delta\mathbf{D}=\mathbf{D}_{\text{obs}}-\mathbf{D}_{\text{model}}$ includes quantities like $D_{V}(z)$, $D_{M}(z)$, and $H(z)$, with $\mathbf{C}_{\text{BAO}}$ as the associated covariance.
Type Ia Supernovae (Pantheon)
The Pantheon sample (Scolnic2018, ) is implemented in MontePython through binned or unbinned distance modulus likelihoods:
$\mathcal{L}_{\text{SN}}\propto\exp\left(-\frac{1}{2}\sum_{j}\frac{\left(\mu_{j}^{\text{obs}}-\mu_{j}^{\text{model}}\right)^{2}}{\sigma_{\mu}(j)^{2}}\right),$ (49)
where $\mu_{j}^{\text{obs}}$ and $\mu_{j}^{\text{model}}$ are the observed and predicted distance moduli, respectively, and $\sigma_{\mu}(j)$ denotes the measurement uncertainties. Full covariance support is also available for high-precision analyses.
CMB Lensing
Lensing likelihoods constrain projected matter distributions and are available via Planck (Planck2018, ) and other CMB experiments. The likelihood is computed from:
$\mathcal{L}_{\text{lensing}}\propto\exp\left(-\frac{1}{2}\Delta\mathbf{d}_{\text{lens}}^{\mathrm{T}}\mathbf{C}_{\text{lens}}^{-1}\Delta\mathbf{d}_{\text{lens}}\right),$ (50)
where $\Delta\mathbf{d}_{\text{lens}}=\mathbf{d}_{\text{obs}}-\mathbf{d}_{\text{theory}}$ corresponds to lensing deflection spectra or convergence maps, and $\mathbf{C}_{\text{lens}}$ is the covariance matrix.
Posterior Sampling and Chi-Squared Evaluation
MontePython performs MCMC sampling using the Metropolis–Hastings algorithm, optionally with adaptive proposals and fast sampling modes. The posterior is defined as:
$\mathcal{P}(\theta|\text{data})\propto\mathcal{L}(\text{data}|\theta)\cdot\pi(\theta),$ (51)
where $\pi(\theta)$ denotes the priors. A typical choice includes flat or logarithmic priors depending on physical parameter ranges.
The corresponding total chi-squared is given by:
$\chi^{2}_{\text{total}}=\sum_{i}\chi^{2}_{i}=\sum_{i}-2\ln\mathcal{L}_{i},$ (52)
where $\chi^{2}_{i}$ corresponds to each likelihood contribution. The best-fit parameters minimize $\chi^{2}_{\text{total}}$, while confidence intervals are derived from the marginalized posterior.
Table IV shows the flat priors for the cosmological parameters based on (Ratra1 ; Aghanim et al. (2018); Aghanim et al. (2018); Di Valentino et al. (2021); Cardona & Sabogal (2023))
@@
$\Omega_{b}h^{2}$ [0.00,0.10][0.00,0.10][ 0.00 , 0.10 ]
$\Omega_{c}h^{2}$ [0.00,0.15][0.00,0.15][ 0.00 , 0.15 ]
$\tau$ [0.00,0.80][0.00,0.80][ 0.00 , 0.80 ]
$n_{s}$ [0.80,1.20][0.80,1.20][ 0.80 , 1.20 ]
$\log[10^{10}A_{s}]$ [1.60,3.90][1.60,3.90][ 1.60 , 3.90 ]
$100\theta_{MC}$ [0.50,10.50][0.50,10.50][ 0.50 , 10.50 ]
${\Omega_{\nu}}$ [0.00,0.009][0.00,0.009][ 0.00 , 0.009 ]
@@ VIII Hubble Tension analysis
${\Omega_{b}h^{2}}$ 0.02226±0.000220.02226$\pm$ 0.000220.02226 ± 0.00022 0.02228±0.000230.02228$\pm$ 0.000230.02228 ± 0.00023 0.02234±0.000200.02234$\pm$ 0.000200.02234 ± 0.00020 0.02236±0.000210.02236$\pm$ 0.000210.02236 ± 0.00021 0.02219±0.000180.02219$\pm$ 0.000180.02219 ± 0.00018
${\Omega_{c}h^{2}}$ 0.1196±0.00370.1196$\pm$ 0.00370.1196 ± 0.0037 0.1194±0.00410.1194$\pm$ 0.00410.1194 ± 0.0041 0.1191±0.00350.1191$\pm$ 0.00350.1191 ± 0.0035 0.1194±0.00350.1194$\pm$ 0.00350.1194 ± 0.0035 0.1187±0.00250.1187$\pm$ 0.00250.1187 ± 0.0025
${100\theta_{MC}}$ 1.0410±0.00061.0410$\pm$ 0.00061.0410 ± 0.0006 1.0408±0.00071.0408$\pm$ 0.00071.0408 ± 0.0007 1.0411±0.00061.0411$\pm$ 0.00061.0411 ± 0.0006 1.04115±0.00051.04115$\pm$ 0.00051.04115 ± 0.0005 1.0408±0.00051.0408$\pm$ 0.00051.0408 ± 0.0005
${\tau}$ 0.059±0.0080.059$\pm$ 0.0080.059 ± 0.008 0.061±0.0090.061$\pm$ 0.0090.061 ± 0.009 0.0551±0.0080.0551$\pm$ 0.0080.0551 ± 0.008 0.055±0.0080.055$\pm$ 0.0080.055 ± 0.008 $0.054^{+0.006}_{-0.007}$
${\mathrm{ln}(10^{10}A_{s})}$ 3.047±0.023.047$\pm$ 0.023.047 ± 0.02 3.045±0.0173.045$\pm$ 0.0173.045 ± 0.017 3.043±0.0193.043$\pm$ 0.0193.043 ± 0.019 3.046±0.0183.046$\pm$ 0.0183.046 ± 0.018 $3.042^{+0.016}_{-0.017}$
${n_{s}}$ 0.971±0.0090.971$\pm$ 0.0090.971 ± 0.009 0.970±0.0090.970$\pm$ 0.0090.970 ± 0.009 0.974±0.0080.974$\pm$ 0.0080.974 ± 0.008 0.964±0.0080.964$\pm$ 0.0080.964 ± 0.008 0.969±0.0070.969$\pm$ 0.0070.969 ± 0.007
@@
CMB + Lensing 69.6±2.869.6$\pm$ 2.869.6 ± 2.8 $0.61\sigma$ $1.26\sigma$
CMB + BAO + Lensing 70.15±1.670.15$\pm$ 1.670.15 ± 1.6 $1.23\sigma$ $1.18\sigma$
CMB + Lensing + CC 70.52±1.9270.52$\pm$ 1.9270.52 ± 1.92 $1.15\sigma$ $1.22\sigma$
CMB + Lensing + Pantheon 70.88±1.9470.88$\pm$ 1.9470.88 ± 1.94 $1.22\sigma$ $1.10\sigma$
CMB + All Data (BAO, CC, Pantheon, Lensing) 70.18±1.7970.18$\pm$ 1.7970.18 ± 1.79 $1.51\sigma$ $1.44\sigma$
@@
${\Omega_{b}h^{2}}$ 0.02230±0.000220.02230$\pm$ 0.000220.02230 ± 0.00022 0.02229±0.000240.02229$\pm$ 0.000240.02229 ± 0.00024 0.02232±0.000210.02232$\pm$ 0.000210.02232 ± 0.00021 0.02233±0.000230.02233$\pm$ 0.000230.02233 ± 0.00023 0.02210±0.000160.02210$\pm$ 0.000160.02210 ± 0.00016
${\Omega_{c}h^{2}}$ 0.1198±0.00360.1198$\pm$ 0.00360.1198 ± 0.0036 0.1194±0.00390.1194$\pm$ 0.00390.1194 ± 0.0039 0.1195±0.00330.1195$\pm$ 0.00330.1195 ± 0.0033 0.1197±0.00340.1197$\pm$ 0.00340.1197 ± 0.0034 0.1186±0.00220.1186$\pm$ 0.00220.1186 ± 0.0022
${100\theta_{MC}}$ 1.0408±0.00061.0408$\pm$ 0.00061.0408 ± 0.0006 1.0409±0.00071.0409$\pm$ 0.00071.0409 ± 0.0007 1.0410±0.00051.0410$\pm$ 0.00051.0410 ± 0.0005 1.04111±0.00061.04111$\pm$ 0.00061.04111 ± 0.0006 1.0406±0.00041.0406$\pm$ 0.00041.0406 ± 0.0004
${\tau}$ 0.058±0.0080.058$\pm$ 0.0080.058 ± 0.008 0.059±0.0090.059$\pm$ 0.0090.059 ± 0.009 0.0550±0.0080.0550$\pm$ 0.0080.0550 ± 0.008 0.055±0.0090.055$\pm$ 0.0090.055 ± 0.009 $0.053^{+0.005}_{-0.007}$
${\mathrm{ln}(10^{10}A_{s})}$ 3.048±0.0193.048$\pm$ 0.0193.048 ± 0.019 3.044±0.0183.044$\pm$ 0.0183.044 ± 0.018 3.045±0.0173.045$\pm$ 0.0173.045 ± 0.017 3.045±0.0173.045$\pm$ 0.0173.045 ± 0.017 $3.041^{+0.015}_{-0.016}$
${n_{s}}$ 0.973±0.0090.973$\pm$ 0.0090.973 ± 0.009 0.971±0.0100.971$\pm$ 0.0100.971 ± 0.010 0.972±0.0080.972$\pm$ 0.0080.972 ± 0.008 0.966±0.0090.966$\pm$ 0.0090.966 ± 0.009 0.968±0.0060.968$\pm$ 0.0060.968 ± 0.006
@@
CMB + Lensing 69.82±3.1269.82$\pm$ 3.1269.82 ± 3.12 $0.74\sigma$ $0.95\sigma$
CMB + BAO + Lensing 70.37±2.3270.37$\pm$ 2.3270.37 ± 2.32 $1.14\sigma$ $1.03\sigma$
CMB + Lensing + CC 70.26±2.3870.26$\pm$ 2.3870.26 ± 2.38 $1.13\sigma$ $0.98\sigma$
CMB + Lensing + Pantheon 70.57±2.5470.57$\pm$ 2.5470.57 ± 2.54 $1.17\sigma$ $0.88\sigma$
CMB + All Data (BAO, CC, Pantheon, Lensing) 70.46±2.0170.46$\pm$ 2.0170.46 ± 2.01 $1.42\sigma$ $1.12\sigma$